How to determine the quadratic equation given the roots or given the sum and product of the roots?
Answer by algebrahouse.com(1659)


Work backwards.

For example:
If x = 2 and x = -3
x - 2 = 0 and x + 3 = 0 {subtracted 2 from first equation and added 3 to second equation}
(x - 2)(x + 3) = 0 {since both equal zero, multiply together and set equal to zero}
x² + x - 6 = 0 {used foil method}

For more help from me, visit: www.algebrahouse.com